Use the five step for problem-solving to answer the following questio.
Angle P and Angle Q are supplementary angles. If, Angle Q is ten less than 4 times Angle P what are the measures of the angles?

Thank you in advance for your help.

Answer by mananth(16949)

let angle P be x
angle Q = 180-x
180-x= 4x-10
5x=190
/5
x=38
Angle P= 38 degrees
angle Q= 180-38= 142 degrees